Harris Company manufactures and sells a single product. A partially completed schedule of the companys total costs and costs per unit over the relevant range of 52,000 to 92,000 units is given below: Required 1. Complete the schedule of the companys total costs and costs per unit as given in the relevant tab below 2. Assume that the company produces and sells 82,000 units during the year at a selling price of $9.84 per unit. Prepare a contribution format income statement for the year. Complete this question by entering your answers in the tabs below Required 1Required 2 Complete the schedule of the companys total costs and costs per unit as given in the relevant tab below. (Round the per unit variable cost and fixed cost to 2 decimal places.) Units Produced and Sold 52,000 72,000 92,000 Total costs Variable costs 192,400 370,000 562,400 Fixed costs Total costs Cost per unit: Variable cost Fixed cost Total cost per unit Required 1 Required 2>Harris Company manufactures and sells a single product. A partially completed schedule of the companys total costs and costs per unit over the relevant range of 52,000 to 92,000 units is given below: Required 1. Complete the schedule of the companys total costs and costs per unit as given in the relevant tab below 2. Assume that the company produces and sells 82,000 units during the year at a selling price of $9.84 per unit. Prepare a contribution format income statement for the year. Complete this question by entering your answers in the tabs below. Required 1Required 2 Assume that the company produces and sells 82,000 units during the year at a selling price of $9.84 per unit. Prepare a contribution format income statement for the year. Harris Company Contribution Format Income Statement KRequired 1 Required 2
Answer #1

Part 1

Units produced and sold
52000 72000 92000
Total costs:
Variable costs 192400 266400 340400
Fixed costs 370000 370000 370000
Total costs 562400 636400 710400
Cost per unit :
Variable cost 3.70 3.70 3.70
Fixed cost 7.12 5.14 4.02
Total cost per unit 10.82 8.84 7.72

Variable cost per unit = 192400/52000 = 3.70

72000*3.70=266400

92000*3.70=340400

370000/52000= 7.12

370000/72000 =5.14

370000/92000 = 4.02

Part 2

Harris company

Contribution format income statement

Sales (82000*9.84) 806880
variable costs (82000*3.70) (303400)
Contribution margin 503480
Fixed costs (370000)
Operating income 133480